A motel clerk counts his $1 and $10 bills at the end of the day. He finds that he has a totalof 60 bills having a combined monetary value of $240. find the number of bills of each denmination that he has
Found 2 solutions by vasumathiest@gmail.com, mananth:
Answer by vasumathiest@gmail.com(2)

Let the $1 bills be x
Then $10 bills would be (60-x)
Total value is ....
1*x + 10(60-x)=x+600-10x=600-9x
but the total value is given and it is $240 so we have ....
600-9x = 240
add 9x on both sides
600-9x+9x=240+9x----->
600 = 240+9x
subtract 240 on both sides
600-240 = 240+9x-240
360 = 9x
divide both sides by 9
360/9 = 9x/9
40 = x
so the $1 bills are 40
and $10 bills are 60-40= 20

Answer by mananth(16946)

A motel clerk counts his $1 and $10 bills at the end of the day. He finds that he has a totalof 60 bills having a combined monetary value of $240. find the number of bills of each denmination that he has
$ 1 bill x numbers
$10 bill 60-x numbers
Value of bills = $240
..
1*x+10*(60-x)=240
x+600-10x=240
-9x=240-600
-9x=-360
/-9
x= 40 $ 1 bills
..
$ 10 bills 20 numbers
...
m.ananth@hotmail.ca